OpenStudy (kjay):

solve the equation for x: 6/x+5 = 7/x+4

hero (hero):

Hint: Cross Multiply

hero (hero):

And use parentheses when posting fractions linearly

OpenStudy (kjay):

here's what I got after cross multiplying: 6x + 24 = 6x + 20 Am I on the right track?

OpenStudy (lgbasallote):

how did you get 6x + 20?

OpenStudy (kjay):

oops, my bad

OpenStudy (lgbasallote):

when you cross multiply it should be 6(x+4) = 7(x+5)

OpenStudy (kjay):

I meant 6x + 35

OpenStudy (kjay):

30*

OpenStudy (lgbasallote):

hmm you should see what i just wrote....

OpenStudy (kjay):

ah, shoot. I just made a small mistake. I accidentally multiplied 6 with both. The equation makes much more sense now.

OpenStudy (lgbasallote):

hehe ;)

OpenStudy (kjay):

I got x = -11 Did I get it correct?

OpenStudy (lgbasallote):

6x + 24 = 7x + 35 7x - 6x = 24 - 35 x = -11

OpenStudy (kjay):

awesome! thank you!